OpenStudy (anonymous):

The graph of f(x) = x2 is shifted 3 units down to obtain the graph of g(x). Which of the following equations best describes g(x)? g(x) = (x + 3)2 g(x) = x2 - 3 g(x) = x2 + 3 g(x) = (x - 3)2 Fan and Medal.. someone please help!!

zepdrix (zepdrix):

Hey Jay c:\[\Large\rm f(x)=x^2\] --------------------------------------------------------- To make a vertical translation we subtract that amount from y,\[\Large\rm v(x)-4=x^2\]adding 4 to the other side,\[\Large\rm v(x)=x^2+4\]That would translate our function 4 units `up`. --------------------------------------------------------- To make a horizontal translation we subtract that amount from x,\[\Large\rm h(x)=(x-5)^2\]In this example, our function would translate 5 to the `right`.

OpenStudy (anonymous):

thanks soo much !! @zepdrix ...

zepdrix (zepdrix):

np c: Able to figure out the correct answer?